# Hestia: Hierarchical Next-Best-View Exploration for Systematic Intelligent Autonomous Data Collection

### v1

@@ Appendix C Qualitative Results
Fig. A1 shows that Hestia’s viewpoints successfully reconstruct the point clouds of real-world scanned objects, while other baselines often miss parts and perform less robustly across various object shapes. Specifically, other baselines miss the teddy bear’s body, the model toy’s face or skirt, the starfish’s arms, the sofa’s front or bottom, the plant’s pot, the statue’s head or stand, the table’s surface, and the durian’s flesh. All these diverse missing parts are captured by Hestia. Fig. A2 shows that Hestia’s viewpoints robustly cover various object shapes, while other baselines often miss finer details or parts underneath. Specifically, other methods miss the wooden stand’s legs, the Lego man’s face or arms, the lamp’s lampshade or neck, the underside of the wooden log, the tree’s leaves, the stairway’s columns or walls, and the cactus’s hat or hands. In contrast, Hestia successfully captures all these diverse and challenging parts. Fig. A3 shows that Hestia’s viewpoints can effectively capture building-like structures, while other methods often miss features such as pillars, roof soffits, or windows.
@@
Figure A4: Qualitative comparison of objects at the four corners. Hestia performs robustly across different object position settings, while other baselines fail to maintain consistent performance across positions.
Fig. A4 shows that Hestia is the only method that achieves consistent performance across different object position settings. These visualization results validate that our proposed core components, including dataset choice, observation design, action space, reward calculation, and learning scheme, form a significant foundation for the tasks and thereby bring a non-marginal impact.
Appendix D Real-World Results
In this section, we present real-world images captured using Hestia operating within the proposed data collection system.
Figure A5: Real-world demonstration of non-shifted and shifted scenes. Hestia can operate in real-world scenarios with three initial viewpoints (red rectangles) and predict the next-best viewpoints for capture (blue rectangles for the first six views), even when the depth camera is unavailable. Point cloud reconstruction results are shown on the left, with black rectangles representing camera poses.
@@ Appendix D Real-World Results
Figure A6: Real-world demonstration of various scenes. Hestia operates in real-world scenarios starting from three initial viewpoints (red rectangles) and predicts the next-best viewpoints for capture (blue rectangles for the first six views), even when the depth camera is unavailable. Point cloud reconstruction results are shown on the left, with black rectangles representing the camera poses.
Fig. A5 demonstrates that Hestia performs well in real-world scenarios, even when the depth camera is unavailable, for both shifted and non-shifted cases. Fig. A6 shows that Hestia robustly handles various real-world object shapes. These results demonstrate that Hestia surpasses prior works, which have not been validated in real-world environments. In addition, Hestia is suitable for use as a viewpoint predictor for real-world applications.
Appendix E Reward Design
@@
$r_{t}=R(s_{t},a_{t})=r_{\text{coverage}}(s_{t},a_{t})+r_{\text{constraint}}(s_{t},a_{t}).$ (A5)
@@
$r_{\text{coverage}}(s_{t},a_{t})=\frac{\sum_{i=1}^{N}\sum_{j=1}^{6}\left(F_{t}^{i,j}-F_{t-1}^{i,j}\right)\cdot M_{\text{col}}}{N\cdot 6}\cdot 0.3,$ (A6)
where $F_{t}^{i,j}$ and $F_{t-1}^{i,j}$ denote the visibility status of the $j$-th face of the $i$-th voxel at time $t$ and $t-1$, respectively. The variable $M_{\text{col}}\in\{0,1\}$ acts as a collision indicator, set to 0 when a collision occurs, thereby nullifying any potential reward for unsafe actions. This reward is computed based on the increment in newly visible voxel faces at the current step. By focusing on the increment rather than the accumulated visibility, the agent is better able to associate rewards with the immediate effects of its actions. To discourage unsafe or invalid decisions, we define a penalty as
$r_{\text{constraint}}(s_{t},a_{t})=\begin{cases}-0.01,&\text{if }r_{\text{coverage}}(s_{t},a_{t})=0,\\ &\text{or }a_{t}[2]>H_{t},\\ &\text{or }a_{t}\in\text{non-free voxels},\\ 0,&\text{otherwise.}\end{cases}$ (A7)
In particular, a negative reward is assigned if the agent fails to reveal any new faces, attempts to move above the maximum allowable flight height $H_{t}$, or selects a viewpoint located within non-free voxels. To ensure a balance between positive and negative rewards, the positive reward is scaled by a factor of 0.30.30.3. This weighting is based on the observation that the maximum face ratio is 1 and each episode ends after 50 steps. As a result, the total possible positive reward is approximately 0.3×1=0.30.3$\times$ 1=0.30.3 × 1 = 0.3, which roughly aligns with the maximum overall value of the negative penalty 0.01×50=0.50.01$\times$ 50=0.50.01 × 50 = 0.5. If the episode ends earlier, such as after 30 steps, this reward structure maintains a perfect balance.
Appendix F Novelty Justification
This section elaborates on the novelty of Hestia.
@@ Appendix F Novelty Justification
Active3D [49]
NeRF-En [54]
ActiveNeRF [50]
NeurAR [56]
UnGuide [51]
ActiveRMAP [52]
MACARONS [57]
NeU-NBV [53]
ScanRL [55]
GenNBV [43]
@@
Table A4: Comparison of learning-based next-best-view methods. Compared to online learning methods, Hestia achieves real-time inference speed. Compared to generalizable methods, Hestia is the only one that performs robustly across different object position settings and demonstrates feasibility in real-world scenarios.
@@
Unlike prior approaches, Hestia systematically addresses the next-best-view task by introducing core components such as dataset selection, observation design, action space formulation, reward computation, and learning schemes. Together, these elements form a comprehensive and unified foundation for the planner (see Sec. H). As shown in Tab. A4, compared to online-learning methods [49, 54, 50, 56, 51, 52, 57], Hestia avoids the need to sample candidate views or perform online optimization. This results in greater flexibility in viewpoint prediction and supports real-time inference. Additionally, in comparison to generalizable methods [53, 55, 43], Hestia is trained on a significantly larger and more diverse dataset, enabling it to generalize robustly to a wide variety of object shapes during testing. Hestia is also the only method that consistently performs well under different object configurations, as demonstrated in Tab. A1, Tab. A2, and Tab. A3. These advantages stem from the key innovations in our design, including treating voxels as cubes rather than points, employing a hierarchical structure to manage the complexity of the action space, and using a greedy learning scheme to mitigate spurious correlations.
Appendix G Spurious Correlation
In this section, we present the spurious correlation caused by future positive rewards in the task.
@@ Appendix G Spurious Correlation
Figure A7: Spurious correlation. In this study, spurious correlation refers to the assignment of future positive rewards to current non-beneficial actions, resulting in suboptimal viewpoint predictions. For instance, the third, thirteenth, and fifteenth viewpoints are empty in the non-greedy design, while enabling the greedy design alleviates this issue.
@@
Spurious correlation has been widely observed across various tasks [62, 63]. In our task, we find that using a large discount factor and future goal rewards can lead to false associations between current actions and their rewards. This creates an illusion for the RL agent that the current action is beneficial, even when there is no information gain (e.g., empty views as shown in Fig. A7) resulting from the current action. Enabling the greedy design mitigates this issue as shown in Fig. A7.
Appendix H Ablation Study
This section validates that Hestia achieves the best performance when all proposed components are combined.
Face Greedy Hier. Avg. CR $\uparrow$ Avg. CD $\downarrow$ #Pa.
Hestia 0.88 0.20 6.2M
0.90 0.17 6.2M
0.95 0.09 6.2M
0.96 0.07 4.9M
Table A5: Ablations. Integrating the proposed ideas with the Objaverse [61, 60] dataset yields top performance with fewer parameters.
We conduct an ablation study (see Tab. A5) on the three concepts of Hestia: using faces as observations, the greedy formulation with a small discount factor and no goal reward, and the hierarchical structure. For simplicity, we keep the input states consistent with Hestia, and for the non-hierarchical structure, we feed the encoded grid information into the policy model to predict 5-DoF viewpoints without feature interpolation. The results show that all these components are essential to improving the performance.
Appendix I Datasets
@@
This section briefly introduces the three main datasets used in the Hestia benchmark: Houses3K [55], Objaverse [61, 60], and OmniObject3D [85].
@@
Figure A8: Sample data from Houses3K [55], Objaverse [60], and OmniObject3D [85]. Houses3K features building shapes with challenging self-occlusions, such as roof soffits. Objaverse includes a diverse range of object shapes. OmniObject3D contains high-quality real-world 3D scans.
I.1 Introduction to Datasets
@@
Houses3K Dataset. Houses3K [55] is designed for next-best-view policy learning. The dataset contains 600 distinct buildings, each rendered with five texture variants, yielding a total of 3,000 FBX models. Many buildings feature challenging self-occlusions, such as roof soffits, that can be fully observed only from bottom-up viewpoints (see Fig. A8). However, because the dataset includes only a single cube-like object category (e.g., buildings), its diversity is limited, which may hinder the ability of next-best-view policies trained on Houses3K to generalize to other structures or everyday objects.
@@
Objaverse Dataset. Objaverse [60] is one of the largest open 3D datasets, containing more than 800,000 shapes across at least 18 high-level categories, including furniture, vehicles, animals, and plants (see Fig. A8). Each category is further divided into several subcategories. The dataset’s scale and diversity make it particularly well suited for foundation model research, especially for 3D generative models. To the best of our knowledge, we are the first to introduce Objaverse for next-best-view policy learning. Its large-scale and diverse object coverage enables training next-best-view policies that perform robustly across a wide range of categories and shapes.
@@
OmniObject3D Dataset. OmniObject3D [85] is a high-quality 3D object dataset collected through real-world scanning, consisting of approximately 6,000 objects across more than 190 categories (see Fig. A8). Unlike synthetic datasets, OmniObject3D captures real-world geometry and texture details using high-resolution 2D and 3D sensors. It provides accurate geometry and realistic material properties, making it commonly used for evaluating real-world transferability in vision tasks such as novel-view synthesis. In this paper, we introduce OmniObject3D for benchmark purposes.
I.2 Dataset Preparation
@@
Figure A9: More diverse and large-scale training set. The chamfer distance measures the discrepancy between the point cloud of the training data and the sphere point cloud. The logarithmic scale of the count represents the number of shapes within each distance range. The right portion displays sample shapes with the same chamfer distance, shown side by side for each dataset. The wider chamfer distance range, higher number of shapes per chunk, and varied shape categories demonstrate that our training data, processed from Objaverse [61, 60], are more comprehensive and large-scale compared to House3K.
@@
We propose using Objaverse [61, 60] as the training dataset to ensure a diverse range of shapes during training (see Fig. A9 and Fig. A10). To achieve this, we filter out large meshes and download the remaining mesh files from Objaverse [61, 60], resulting in a dataset comprising 120,000 shapes. For each shape, we generate the occupancy grid and point cloud using Open3D [86]. To remove invisible voxels and points, we perform a breadth-first search (BFS) starting from external free voxels, retaining only reachable occupancy voxels and points as the ground truth. The visible faces of each voxel are identified by examining the occupancy states of neighboring voxels. We apply PCA [87] to reduce the point cloud to three components and use k-means clustering [88] to group the 120,000 shapes into 30,100 clusters. The cluster centers of 30,000 clusters are designated as training data, while the remaining cluster centers are used for testing. The same procedure is used to create 256 training samples and 100 test samples from the House3K dataset [55] for our benchmark.
@@
Figure A10: Dataset distribution. Our training dataset, processed from Objaverse [61, 60], includes a wide range of categories and is not limited to cubic-like shapes (e.g., buildings).
@@
Our processed training set is two orders of magnitude larger than those used in previous studies [53, 43, 55] and includes at least 18 more categories than prior datasets [43, 55]. As for OmniObject3D [85], due to limited storage space, we randomly select one shape per category for evaluation, resulting in approximately 200 test samples.
Appendix J Real-World Data Collection System
@@
This section includes the setup and pseudo code of the proposed real-world data collection system. By integrating Hestia into the proposed real-world data collection system, Hestia demonstrates its feasibility under practical scenarios, surpassing prior approaches.
J.1 Real-World System Setup
The environment size of our scenarios (e.g., an opera house) is approximately 2.6 m $\times$ 2.6 m $\times$ 2 m. To prevent the drone from exceeding the HTC Lighthouse base station range, the maximum height $H_{t}$ is restricted to 1.5 m. Additionally, in the nearest collision-free voxel module, voxels below 0.4m are marked as occupied to avoid potential counterforces between the floor and the drone’s quadrotor. In this system, we deploy the DJI Mini 3 Pro as the primary aircraft model.
We integrate the DJI Mobile SDK v5 to enable remote control and command transmission from the base station to the UAV. This software development kit provides developers with comprehensive control capabilities over the UAV. The SDK is embedded in an Android application package, where we develop a custom application capable of broadcasting aircraft data via the User Datagram Protocol (UDP) wireless network protocol. The broadcasted data is captured using a Raspberry Pi 4, which runs a ROS 2 node designed to receive UDP packets and convert them into ROS 2-compatible messages. On the same Raspberry Pi 4, we implement an adaptive trajectory planning technique. This method evaluates a pre-generated library of feasible offline trajectories, allowing the UAV to navigate autonomously by selecting the most appropriate path based on real-time conditions. Integrating these components ensures a reliable flow of data and commands, enabling efficient autonomous navigation for the UAV. We utilize the Crazyflie v2.1, a nano UAV equipped with a Lighthouse Positioning Deck, to achieve precise localization within the experimental environment. To integrate its capabilities with the primary aircraft, we remove the propellers and motors of the Crazyflie and securely mount it on top of the DJI Mini 3 Pro. This setup enables the Crazyflie to serve as a localization beacon, providing accurate positional data for the main UAV within the tracking range of the Lighthouse base stations. The positioning deck on the Crazyflie captures localization data using infrared signals from the Lighthouse system. This data is transmitted wirelessly via the Crazyradio 2.0 module, which connects to a base station. The base station, running the Crazyswarm 2.0 package on the ROS 2 framework, processes and publishes the localization data in real time. This setup facilitates autonomous navigation and precise positioning of the main UAV by continuously updating its coordinates within the experimental space.
To ensure reproducibility, our system is constructed entirely from publicly available, low-cost hardware, and we will release the accompanying code to support open research. The total cost of the additional components, including the Crazyflie v2.1 (approximately \$200), the Lighthouse Positioning Deck (around \$100), and the Crazyradio 2.0 module (about \$50), is approximately \$350. When combined with the DJI Mini 3 Pro, which costs around \$800, the total system cost remains significantly lower than that of conventional localization solutions. This cost-effective design, combined with open-source software such as ROS 2 and the Crazyswarm package, provides a reliable and accessible prototype for UAV-based data collection.
J.2 Real-World System Processes
@@
11:  Receive_Image(i)$\triangleright$ Transmit image (Alg. 4)
@@
15:  Set_$\mathcal{W}$$\triangleright$ Move to NBV (Alg. 2)
@@
The flowchart of the system in the right part of Fig. 4, highlights four main sub-processes: Set Waypoints Process, Capture Image Process, Receive Image Process, and NBV Prediction Process. Each sub-process is represented by distinct colors in the diagram and described as follows:
@@
    2:while $\mathcal{D}_{\text{drone}}.\text{loc}\not\approx\mathcal{D}_{\text{pi}}.\text{x}$ $\triangleright$ Wait until location matches NBV
@@
[…]
@@
Appendix K Training and Testing Details

### v2

@@ 6 Acknowledgements
In this section, we present additional qualitative results for OmniObject3D [wu2023omniobject3d], Objaverse [Deitke_2023_CVPR], and Houses3K [peralta2020next] in Figs. S2, S3, S4, 3 and S6. Fig. S2 shows that Hestia’s viewpoints successfully reconstruct the point clouds of real-world scanned objects, while other baselines often miss parts and perform less robustly across various object shapes. Specifically, other baselines miss the starfish’s arms, the sofa’s front or bottom, the plant’s pot, the statue’s head or stand, the table’s surface, and the durian’s flesh. All these diverse missing parts are captured by Hestia. Fig. S3 shows that Hestia’s viewpoints robustly cover various object shapes, while other baselines often miss finer details or parts underneath. Specifically, other methods miss the wooden stand’s legs, the Lego man’s face or arms, the lamp’s lampshade or neck, the underside of the wooden log, and the tree’s leaves. In contrast, Hestia successfully captures all these diverse and challenging parts. Fig. S4 shows that Hestia’s viewpoints can effectively capture building-like structures, while other methods often miss features such as pillars, roof soffits, or windows. Fig. S5 shows that Hestia can capture the complex scene well. Fig. S6 shows that Hestia is the only method that achieves consistent performance across different object position settings. These visualization results validate that our proposed core components, including dataset choice, observation design, action space, reward calculation, and learning scheme, form a significant foundation for the tasks and thereby bring a non-marginal impact.
Figure S2: Qualitative comparison on OmniObject3D [wu2023omniobject3d]. Hestia’s viewpoints reconstruct the point clouds of real-world scanned objects accurately, while other baselines exhibit less robustness across various object shapes, often missing parts in the reconstructed point clouds.
Figure S3: Qualitative comparison on Objaverse [Deitke_2023_CVPR]. Hestia’s viewpoints successfully reconstruct diverse and complex object shapes, while other baselines exhibit less robustness, often missing self-occluded regions or parts that require bottom-up viewpoints.
Figure S4: Qualitative comparison on Houses3K [peralta2020next]. Compared to the baselines, the point clouds reconstructed from the depth maps collected by Hestia capture finer details, such as roof soffits, pillars, and windows, particularly in self-occluded areas.
Figure S5: Reconstruction on a complex scene. Hestia captures the scene well.
@@
Figure S6: Qualitative comparison of objects at the four corners. Hestia performs robustly across different object position settings, while other baselines fail to maintain consistent performance across positions.
S4 Real-World Results
In this section, we present real-world images captured using Hestia operating within the drone system (see Sec. S9). Fig. 5 demonstrates that Hestia performs well in real-world object-centric scenes, even when the depth camera is unavailable, for both shifted and non-shifted cases. Notably, without a depth camera to synchronize the virtual and real world, we manually set up three viewpoints, which are deliberately placed close to each other (e.g., the red boxes in Fig. 5). In addition, it is reasonable that some next-best viewpoints appear similar because we use a multi-view depth predictor [duisterhof2024mast3r, wang2024dust3r] to convert RGB images into depth maps. Therefore, it is common for certain viewpoints to overlap in order to obtain depth and update the input state. Fig. S7 shows that Hestia robustly handles various real-world object shapes. These results demonstrate that Hestia surpasses the prior works [guedon2023macarons, peralta2020next, chen2024gennbv, jin2023neu, lin2022active, sunderhauf2023density, pan2022activenerf, ran2023neurar, zhan2022activermap], which have not been validated in real-world environments. In addition, Hestia is suitable for use as a viewpoint predictor for real-world applications.
@@
Figure S7: Real-world demonstration of various object-centric scenes. Hestia operates in real-world object-centric scenes starting from three initial viewpoints (red rectangles) and predicts the next-best viewpoints for capture (blue rectangles for the first six views), even when the depth camera is unavailable. Point cloud reconstruction results are shown on the left, with black rectangles representing the camera poses.
S5 Reward Design
@@
$r_{t}=R(s_{t},a_{t})=r_{\text{coverage}}(s_{t},a_{t})+r_{\text{constraint}}(s_{t},a_{t}).$ (S5)
@@
$r_{\text{coverage}}(s_{t},a_{t})$ $=\frac{\sum_{i=1}^{N}\sum_{j=1}^{6}\left(F_{t}^{i,j}-F_{t-1}^{i,j}\right)\cdot M_{\text{col}}}{N\cdot 6}$
$\quad\cdot 0.3$ (S6)
where $F_{t}^{i,j}$ and $F_{t-1}^{i,j}$ denote the visibility status of the $j$-th face of the $i$-th voxel at time $t$ and t1t-1, respectively. The variable $M_{\text{col}}\in\{0,1\}$ acts as a collision indicator, set to 0 when a collision occurs, thereby nullifying any potential reward for unsafe actions. This reward is computed based on the increment in newly visible voxel faces at the current step. By focusing on the increment rather than the accumulated visibility, the agent is better able to associate rewards with the immediate effects of its actions. To discourage unsafe or invalid decisions, we define a penalty as
$r_{\text{constraint}}(s_{t},a_{t})=\begin{cases}-0.01,&\text{if }r_{\text{coverage}}(s_{t},a_{t})=0,\\ &\text{or }a_{t}[2]>H_{t},\\ &\text{or }a_{t}\in\text{non-free voxels},\\ 0,&\text{otherwise.}\end{cases}$ (S7)
In particular, a negative reward is assigned if the agent fails to reveal any new faces, attempts to move above the maximum allowable flight height $H_{t}$, or selects a viewpoint located within non-free voxels. To ensure a balance between positive and negative rewards, the positive reward is scaled by a factor of $0.3$. This weighting is based on the observation that the maximum face ratio is 1 and each episode ends after 50 steps. As a result, the total possible positive reward is approximately $0.3\times 1=0.3$, which roughly aligns with the maximum overall value of the negative penalty $0.01\times 50=0.5$. If the episode ends earlier, such as after 30 steps, this reward structure maintains a perfect balance.
S6 Novelty Justification
@@
Active3D [lin2022active]
NeRF-En [sunderhauf2023density]
ActiveNeRF [pan2022activenerf]
NeurAR [ran2023neurar]
EfficientView [zeng2022efficient]
UnGuide [lee2022uncertainty]
SEER [tao2022seer]
ActiveRMAP [zhan2022activermap]
MACARONS [guedon2023macarons]
NeU-NBV [jin2023neu]
ScanRL [peralta2020next]
GenNBV [chen2024gennbv]
@@
Table S4: Comparison of learning-based next-best-view methods. Compared to online learning methods, Hestia achieves real-time inference speed. Compared to generalizable methods that predict five-degree-of-freedom viewpoints or assume a drone as the agent, Hestia exhibits robustness across different object position settings and demonstrates feasibility in real-world object-centric scenes.
@@
This section elaborates on the novelty of Hestia. Hestia is a generalizable next-best-view planner that can predict five-degree-of-freedom viewpoints and model a drone as an agent. Therefore, we mainly focus on comparing methods that also predict five-degree-of-freedom viewpoints or assume a drone as an agent. Unlike prior approaches, Hestia systematically addresses the next-best-view task by introducing core components such as dataset selection, observation design, action space formulation, reward computation, and learning schemes. Together, these elements form a comprehensive and unified foundation for the planner (see Sec. 4.7). As shown in Tab. S4, compared to online-learning methods [lin2022active, sunderhauf2023density, pan2022activenerf, ran2023neurar, lee2022uncertainty, zhan2022activermap, guedon2023macarons], Hestia avoids the need to sample candidate views or perform online optimization. This results in greater flexibility in viewpoint prediction and supports real-time inference. Additionally, in comparison to generalizable methods [jin2023neu, peralta2020next, chen2024gennbv], Hestia is trained on a significantly larger and more diverse dataset, enabling it to generalize robustly to a wide variety of object shapes during testing. Hestia is also the only method that consistently performs well under different object configurations, as demonstrated in Tabs. S1, S2 and S3. These advantages stem from the key innovations in our design, including treating voxels as cubes rather than points, employing a hierarchical structure to manage the complexity of the action space, and using a greedy learning scheme to mitigate spurious correlations. Notably, the purpose of Hestia’s hierarchical structure is to address the high-dimensional continuous action search space problem in reinforcement learning-based generalizable next-best-view planning, which is fundamentally different from traditional methods that use hierarchical structures to move along frontiers. One of the most recent works [chen2024gennbv] still lacks the designs we propose.
S7 Spurious Correlation
@@
Figure S8: Spurious correlation. In this study, spurious correlation refers to the assignment of future positive rewards to current non-beneficial actions, resulting in suboptimal viewpoint predictions. For instance, the third, thirteenth, and fifteenth viewpoints are empty in the non-greedy design, while enabling the close-greedy design alleviates this issue.
@@
In this section, we present the spurious correlation caused by future positive rewards in the task. Spurious correlation has been widely observed across various tasks [interpretation1971spurious, kim2024discovering]. In our task, we find that using a large discount factor and future goal rewards can lead to false associations between current actions and their rewards. This creates an illusion for the reinforcement learning agent that the current action is beneficial, even when there is no information gain (e.g., empty views as shown in Fig. S8) resulting from the current action. Enabling the close-greedy design mitigates this issue as shown in Fig. S8.
S8 Datasets
@@
This section briefly introduces the three main datasets used in the Hestia benchmark: Houses3K [peralta2020next], Objaverse [deitke2024objaverse, Deitke_2023_CVPR], and OmniObject3D [wu2023omniobject3d].
@@
Figure S9: Sample data from Houses3K [peralta2020next], Objaverse [Deitke_2023_CVPR], and OmniObject3D [wu2023omniobject3d]. Houses3K features building shapes with challenging self-occlusions, such as roof soffits. Objaverse includes a diverse range of object shapes. OmniObject3D contains high-quality real-world 3D scans.
S8.1 Introduction to Datasets
@@
Houses3K Dataset. Houses3K [peralta2020next] is designed for next-best-view policy learning. The dataset contains 600 distinct buildings, each rendered with five texture variants, yielding a total of 3,000 FBX models. Many buildings feature challenging self-occlusions, such as roof soffits, that can be fully observed only from bottom-up viewpoints (see Fig. S9). However, because the dataset includes only a single cube-like object category (e.g., buildings), its diversity is limited, which may hinder the ability of next-best-view policies trained on Houses3K to generalize to other structures or everyday objects.
@@
Objaverse Dataset. Objaverse [Deitke_2023_CVPR] is one of the largest open 3D datasets, containing more than 800,000 shapes across at least 18 high-level categories, including furniture, vehicles, animals, and plants (see Fig. S9). Each category is further divided into several subcategories. The dataset’s scale and diversity make it particularly well-suited for foundation model research, especially for 3D generative models. To the best of our knowledge, we are the first to introduce Objaverse for next-best-view policy learning. Its large-scale and diverse object coverage enables training next-best-view policies that perform robustly across a wide range of categories and shapes.
@@
OmniObject3D Dataset. OmniObject3D [wu2023omniobject3d] is a high-quality 3D object dataset collected through real-world scanning, consisting of approximately 6,000 objects across more than 190 categories (see Fig. S9). Unlike synthetic datasets, OmniObject3D captures real-world geometry and texture details using high-resolution 2D and 3D sensors. It provides accurate geometry and realistic material properties, making it commonly used for evaluating real-world transferability in vision tasks such as novel-view synthesis. In this paper, we introduce OmniObject3D for benchmark purposes.
S8.2 Dataset Preparation
@@
Figure S10: More diverse and large-scale training set. The chamfer distance measures the discrepancy between the point cloud of the training data and the sphere point cloud. The logarithmic scale of the count represents the number of shapes within each distance range. The right portion displays sample shapes with the same chamfer distance, shown side by side for each dataset. The wider chamfer distance range, higher number of shapes per chunk, and varied shape categories demonstrate that our training data, processed from Objaverse [deitke2024objaverse, Deitke_2023_CVPR], are more comprehensive and large-scale compared to Houses3K.
@@
We propose using Objaverse [deitke2024objaverse, Deitke_2023_CVPR] as the training dataset to ensure a diverse range of shapes during training (see Figs. S10 and S11). To achieve this, we filter out large meshes and download the remaining mesh files from Objaverse [deitke2024objaverse, Deitke_2023_CVPR], resulting in a dataset comprising 120,000 shapes. For each shape, we generate the occupancy grid and point cloud using Open3D [zhou2018open3d]. To remove invisible voxels and points, we perform a breadth-first search (BFS) starting from external free voxels, retaining only reachable occupancy voxels and points as the ground truth. The visible faces of each voxel are identified by examining the occupancy states of neighboring voxels. We apply PCA [abdi2010principal] to reduce the point cloud to three components and use k-means clustering [macqueen1967some] to group the 120,000 shapes into 30,100 clusters. The cluster centers of 30,000 clusters are designated as training data, while the remaining cluster centers are used for testing. The same procedure is used to create 256 training samples and 100 test samples from the Houses3K dataset [peralta2020next] for our benchmark.
@@
Figure S11: Dataset distribution. Our training dataset, processed from Objaverse [deitke2024objaverse, Deitke_2023_CVPR], includes a wide range of categories and is not limited to cubic-like shapes (e.g., buildings).
@@
Our processed training set is two orders of magnitude larger than those used in previous studies [jin2023neu, chen2024gennbv, peralta2020next] and includes at least 18 more categories than prior datasets [chen2024gennbv, peralta2020next]. As for OmniObject3D [wu2023omniobject3d], due to limited storage space, we randomly select one shape per category for evaluation, resulting in approximately 200 test samples.
Figure S12: Real-world drone system. Hestia is a generalizable next-best-view planner that is feasible for real-world deployment. Please refer to our demonstration video for further details.
Figure S13: Overview of the real-world drone system. The system uses a drone with an RGB camera for data capture, Lighthouse base stations and Crazyflie for localization, and a Wifi router for wireless communication.
S9 Real-World Drone System
@@
This section includes the setup and pseudo code of the real-world drone system we used. By integrating Hestia into the real-world drone system (see Fig. S12), Hestia demonstrates its feasibility under practical scenarios.
@@
10:  $i\leftarrow\texttt{Capture\_Image()}\triangleright$ Capture an image (Alg. 3)
[…]
@@
14:  $\mathcal{W}\leftarrow\texttt{NBV\_Prediction()}\triangleright$ Predict NBV (Alg. 5)
@@
16:  $i\leftarrow\texttt{Capture\_Image()}\triangleright$ Capture an image (Alg. 3)
@@
S9.1 Real-World System Overview
To demonstrate Hestia’s feasibility in a real-world environment, we use a real-world system (see Fig. S13), where a drone equipped with an RGB camera moves to the next-best viewpoint predicted by Hestia to capture images of an object. The system uses four HTC Lighthouse base stations and a Crazyflie 2.1 for localization and transmits images to the ground control station via wireless communication. MASt3R [duisterhof2024mast3r] is integrated to convert RGB images into pointmaps (e.g., depth images), and three initial viewpoints are set for real-world and virtual-world synchronization [umeyama1991least]. Fig. S13 and Alg. 1 illustrate four key processes of the system. Specifically, the drone captures images at three initial viewpoints, where the set waypoints process navigates the drone using a heuristic trajectory planner based on prior knowledge of the environment. Then, the capture image process commands image capture, and the receive image process transmits the image to the ground station. After capturing the initial viewpoints, the nbv prediction process predicts the next-best viewpoint based on the collected data. Four processes repeat until sufficient data is collected. For more details, please refer to Sec. S9.3.
S9.2 Real-World System Setup
The environment size of our object-centric scenes (e.g., an opera house) is approximately 2.6 m $\times$ 2.6 m $\times$ 2 m. To prevent the drone from exceeding the HTC Lighthouse base station range, the maximum height $H_{t}$ is restricted to 1.5 m. Additionally, in the nearest collision-free voxel module, voxels below 0.4m are marked as occupied to avoid potential counterforces between the floor and the drone’s quadrotor. In this system, we deploy the DJI Mini 3 Pro as the primary aircraft model.
We integrate the DJI Mobile SDK v5 to enable remote control and command transmission from the base station to the UAV. This software development kit provides developers with comprehensive control capabilities over the UAV. The SDK is embedded in an Android application package, where we develop a custom application capable of broadcasting aircraft data via the User Datagram Protocol (UDP) wireless network protocol. The broadcast data is captured using a Raspberry Pi 4, which runs a ROS 2 node designed to receive UDP packets and convert them into ROS 2-compatible messages. On the same Raspberry Pi 4, we implement an adaptive trajectory planning technique. This method evaluates a pre-generated library of feasible offline trajectories, allowing the UAV to navigate autonomously by selecting the most appropriate path based on real-time conditions. Integrating these components ensures a reliable flow of data and commands, enabling efficient autonomous navigation for the UAV. We utilize the Crazyflie v2.1, a nano UAV equipped with a Lighthouse Positioning Deck, to achieve precise localization within the experimental environment. To integrate its capabilities with the primary aircraft, we remove the propellers and motors of the Crazyflie and securely mount it on top of the DJI Mini 3 Pro. This setup enables the Crazyflie to serve as a localization beacon, providing accurate positional data for the main UAV within the tracking range of the Lighthouse base stations. The positioning deck on the Crazyflie captures localization data using infrared signals from the Lighthouse system. This data is transmitted wirelessly via the Crazyradio 2.0 module, which connects to a base station. The base station, running the Crazyswarm 2.0 package on the ROS 2 framework, processes and publishes the localization data in real time. This setup facilitates autonomous navigation and precise positioning of the main UAV by continuously updating its coordinates within the experimental space.
The system is constructed entirely from publicly available, low-cost hardware. The total cost of the additional components, including the Crazyflie v2.1 (approximately \$200), the Lighthouse Positioning Deck (around \$100), and the Crazyradio 2.0 module (about \$50), is approximately \$350. When combined with the DJI Mini 3 Pro, which costs around \$800, the total system cost remains significantly lower than that of conventional localization solutions. This cost-effective design, combined with open-source software such as ROS 2 and the Crazyswarm package, provides a reliable and accessible prototype for UAV-based data collection.
S9.3 Real-World System Processes
The flowchart of the system in the right part of Fig. S13, highlights four main sub-processes: Set Waypoints Process, Capture Image Process, Receive Image Process, and NBV Prediction Process. Each sub-process is represented by distinct colors in the diagram and described as follows:
@@
[…]
@@
    2:[$\mathcal{I}\leftarrow[i_{1},\dots,i_{n}]\triangleright$ Load images
@@
S10 Training and Testing Details
Figure S14: Failure cases of Hestia. Hestia may occasionally fail to capture finer 3D structures, highly self-occluded parts, nearly vertical bottom-up views, and small details on coarse object surfaces.